i would like work shown please,using the substitution method
x=-3-y
5x-3y=1

Answer by TutorDelphia(193)

x=-3-y
5x-3y=1
We already have this solved for x, so lets plug in for x
5(-3-y)-3y=1
distribute the 5
-15-5y-3y=1
combine like terms
-15-8y=1
add 15 to both sides
-8y=16
divide both sides by -8
y=-2
And plug in two for y
x=-3-(-2)=-3+2
x=-1